one second after projection, a stone moves at an angle of 45 degree with horizontal. Two second after projection it moves horizontally. Its angle of projection is?

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
after 2 second it becomes horizontal so time of flight=4sec
2*v0*sin(theta)/g=4 =>v0*sin(theta) =2g
vertical velocity let after one second v
vsin45=v0*sin(theta)-g so vcos45=v0*sin(theta)-g
horizontal component remains constant so
v0cos(theta)=v0*sin(theta)-g
=>v0*cos(theta)=v0*sin(theta)/2
=>tan(theta)=2
theta =tan^-1(2)


sher mohammad
askiitians faculty